OpenStudy (anonymous):

The slopes of perpendicular lines are ______ ______ of each other, which means their product is -1. A. negative reciprocals B. positive reciprocals C. fractional reciprocals

OpenStudy (anonymous):

@UsukiDoll @butterflydreamer

OpenStudy (anonymous):

it's okay lol i need help though :(

OpenStudy (butterflydreamer):

what do you think it is?

OpenStudy (anonymous):

C ?

OpenStudy (butterflydreamer):

not quiteee...

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

i dont think its b

OpenStudy (butterflydreamer):

okay so let's saayy... we have perpendicular lines and we are told \[m_1 = 2 , m_2 = -\frac{ 1 }{ 2 }\] focussing on m_1 = 2 what is the reciprocal of 2?

OpenStudy (anonymous):

I dont know nothing about reciprocals 1 ?

OpenStudy (butterflydreamer):

reciprocal means to flip the number basically.. m = 2 = 2/1 right? the reciprocal will be 1/2 because we flipped 2/1 makes sense? but notice how m_2 = -1/2 ? isn't this negative of the reciprocal of m_1 = 2?

OpenStudy (anonymous):

yes

OpenStudy (butterflydreamer):

yep.. so, what would the answer be?

OpenStudy (anonymous):

negative

OpenStudy (butterflydreamer):

yuupppp

OpenStudy (anonymous):

thanks bae (:

OpenStudy (butterflydreamer):

yw.